Two brothers, Andile and Beki, want to join a gym for one year.
Andile decides to join Gym Aerolux and Beki decides to join
Gym Bodyworx. The monthly fee for Aerolux is R125 with no
registration fee and the monthly fee for Bodyworx is R100 with
R250 registration fee.
1
Expert's answer
2020-07-16T19:10:51-0400

Let's calculate how much each brother will pay for the year in total.

IThere are 12 months in a year, so:


Gym Aerolux: "125 * 12 = 1500"

Gym Bodyworx: "100*12+250=1200+250=1450"


So it will be R50 cheaper to join Gym Bodyworx.